solve the system below of linear equations. Verify your solution.
4x + 3y = 13
5x + y -8 = 0
Solve for X and Y.

Answer by suruman(21)

4x + 3y = 13 ....(a)
5x + y - 8 = 0 ....(b)
Multiplying (b) by -3,
-15x -3y +24 = 0 ...(c)
Adding (a) with (c),
-11x +24 = 13
11x = 11
x = 1
Substituting value of x in (a),
4 + 3y = 13
3y = 9
y = 3
Thus the solution is (1,3).
Substituting the above values of x and y in equation (a),
4*1 + 3*3 = 4 + 9 = 13
Substituting the above values of x and y in equation (b),
5*1 + 3 -8 = 5 + 3 - 8 = 8 - 8 =0
Thus, (1,3) is the solution pair for the given set of equations.